The sum of four consecutive integers is -42. Find the absolute value of the largest integer.
Found 3 solutions by CubeyThePenguin, MathTherapy, ikleyn:
Answer by CubeyThePenguin(3113)

4 consecutive integers: x, (x+1), (x+2), (x+3)

x + (x+1) + (x+2) + (x+3) = -42
4x + 6 = -42
4x = -48
x = -8

absolute value of largest integer = |x + 3| = |-8 + 3| = 5

Answer by MathTherapy(10550)


The sum of four consecutive integers is -42. Find the absolute value of the largest integer.
That person's answer is WRONG, WRONG, WRONG!! 


Answer by ikleyn(52767)

.

4 consecutive integers: x, (x+1), (x+2), (x+3)

x + (x+1) + (x+2) + (x+3) = -42
4x + 6 = -42
4x = -48
x = -12

The numbers are -12, -11, -10, -9.

absolute value of largest integer is   |-9| = 9.              ANSWER

------------

Solved (correctly).